Solve by substitution:
-4x - y = 4
x = y + 9

Answer by amarjeeth123(569)

-4x - y = 4.........equation 1
x = y + 9...........equation 2
From 2 we get y=x-9
Substituting in 1 we get,
-4x-x+9=4
-5x=4-9
-5x=-5
x=1.
y=-8.